• An explorer walks 22km due North and then 45 degrees south of east for 47km.what’s the resultant displacement from the origin?
1
Expert's answer
2022-03-16T05:40:20-0400

resultant displacement ="\sqrt{47^2+22^2-2\times 47\times22 cos45^0}=\sqrt{1230.703177}=35.08km"